The area of a rectangle is 48 square centimeters. If the length of the rectangle is 8 centimeters longer than the width, how wide is the rectangle?

Answer by CharlesG2(834)

The area of a rectangle is 48 square centimeters. If the length of the rectangle is 8 centimeters longer than the width, how wide is the rectangle?
A area = L length * W width
48=(W+8)*W
48=W^2 + 8W
0=W^2 + 8W - 48
0=(W+12)(W-4)
W=-12 or W=4, well width can not be negative so, W=4
check:
48=(W+8)*W
48=(4+8)*4
48=12*4
48=48